# MIRACLE II: Unveiling the multi-phase gas interplay in the circumnuclear region of NGC 1365 via multi-cloud modeling

### v2

@@ C.1 JWST/MIRI MRS emission line fitting
This appendix provides further details on the customized python routine we employed to analyze the MIRI MRS datacubes. We followed a three-step procedure: first, we applied spatial smoothing to enhance the SNR; second, we subtracted the local continuum around the brightest emission lines; finally, we performed spaxel-by-spaxel Gaussian fitting of the residual emission. First, we applied a Gaussian spatial smoothing with a kernel size of 1 pixel to every spatial plane of the datacube. This step improves the SNR of faint extended features, at the cost of a moderate degradation in spatial resolution. Specifically, the effective PSF is broadened by approximately 40 across all wavelengths, due to the fact that the pixel size is comparable to or slightly smaller than the intrinsic PSF at each wavelength. Nevertheless, the resulting resolution remains sufficient for our spatially resolved analysis, as it still allows us to distinguish the relevant morphological and kinematic structures in the data.
@@
(1) ISO Spectrometer Data Centre.
@@
In the mid-IR, dust attenuation is much lower than at optical wavelengths, but it does vary with wavelength and affect emission lines differently, potentially impacting specific line ratios. Unfortunately, in our data the SNR of the Pf and Hu (i.e. H I (6-5) at 7.46 m and H I (7-6) at 12.37 m, respectively) is too low to derive a reliable extinction map directly from the mid-IR spectra. To test the potential impact of dust attenuation, we applied an extinction correction to the mid-IR emission lines using the Gordon et al. (2023) attenuation curve and adopting the maximum value (3) derived from our best-fit photoionization model (see Section 3.5.3). However, we find that the resulting variations in the mid-IR line ratios are within the errors, including those involving emission lines at distant wavelengths (e.g., [Ne V]/[Ne V], see Section 3.4).
@@ C.2 MUSE WFM emission line fitting
MUSE data were analyzed as already described in previous works of the MAGNUM program (Venturi et al. 2018; Mingozzi et al. 2019). Briefly, after a preliminary Voronoi tessellation (Cappellari & Copin 2003) to achieve an average SNR per wavelength channel of at least 70 on the continuum, we applied the penalized PiXel-Fitting (pPXF, Cappellari & Emsellem 2004) code to measure the optical emission lines listed in Table 3. Using the penalized PiXel-Fitting (pPXF, Cappellari & Emsellem 2004) code, we fitted simultaneously the continuum and gas emission in the wavelength range 4800-9200 Å, convolving a linear combination of stellar templates (E-MILES; for details see Vazdekis et al. 2016) with a Gaussian velocity distribution. Each emission line was fitted twice, using one and two Gaussian components, and the KS test was applied spaxel-by-spaxel to determine the minimum number of components required to adequately reproduce the line profiles (see Appendix C.1). Given that NGC 1365 hosts a Seyfert 1.8 nucleus, we included an additional spectral component when fitting the bins within a radius of 10 spaxels from the nuclear position to account for the spatially unresolved broad-line region (BLR). The BLR model consists of two independent components: a broad (¿1000 km s-1) Gaussian component for the broad Balmer lines (i.e. H and H) and a template of broad [Fe II] emission features by Boroson & Green (1992). The nuclear model spectrum was obtained by fitting the integrated spectrum of the central spaxels and, in the subsequent spaxel-by-spaxel fitting, it was kept fixed in shape and velocity while being allowed to scale in flux. We then subtracted the stellar and nuclear emission from the original unbinned data cube by rescaling the fitted continuum - assumed constant within each Voronoi bin - to match the observed continuum flux in each spaxel, before performing the subtraction.
@@
Finally, we corrected the optical emission-line fluxes for dust attenuation using the Balmer decrement (i.e., H/H), assuming a case-B hydrogen recombination value of 2.86 for this ratio (for an electron temperature and density of Te = K and ne = cm-3, respectively; Osterbrock & Ferland 2006) and adopting the Gordon et al. (2023) attenuation law. The results, presented in Appendix D, are consistent with previous estimates by Venturi et al. (2018) and Mingozzi et al. (2019). As shown in Fig. 12, the dust attenuation peaks at AV = 3.7 in the circumnuclear ring, where young stellar clusters are embedded (Liu et al. 2023; Schinnerer et al. 2023). High attenuation values are also found at the nuclear position, while we do not find any correlation of AV with the outflow morphology, indicative of its low dust content.
@@ Appendix D Extinction and electron density from MUSE
In this appendix, we present the spatially resolved maps of visual extinction and electron density in the circumnuclear region of NGC 1365, derived from the MUSE data and zoomed-in FoV matching the MIRI one. In the top panel of Fig. 12, we show the visual extinction computed from the Balmer decrement (H/H), adopting the extinction law by Gordon et al. (2023). High visual extinction values of A 3.7 mag are found at the nuclear position and in the southeastern part of the FoV, corresponding to the circumnuclear ring, as also observed by Venturi et al. (2018) and Mingozzi et al. (2019).
@@ Appendix E Density comparison with IRS observations
In this appendix, we compare our spatially resolved JWST/MIRI density measurements with those obtained from integrated spectra by Dudik et al. (2007). Dudik et al. (2007) used the Infrared Spectrograph (IRS) on board Spitzer to compute the [Ne V] ratio in a sample of local AGN. They found that Type 2 AGN show higher line ratios with respect to Type 1, with all Type 2 sources lying above the theoretical LDL threshold, implying log(ne/cm-3)2. In particular, for NGC 1365 they computed a [Ne V]/[Ne V] ratio of 2.40.6, which is larger than our estimation of 1.00.2 obtained as the ratio between the two integrated fluxes in the MIRI FoV. They argued that the [Ne V] emitting region likely originates inside the nuclear torus and that differential infrared extinction due to dust in the obscuring torus may be responsible for the ratio above the LDL. Exploiting the spatially resolved [Ne V] emission in our data and the density values in the nuclear region, we can rule out this interpretation. However, given the presence of some LDL regions in Fig. 6, it is possible that their finding is the result of spatial integration over a wide slit, thus including such regions in the measurement. A more plausible reason is an observational effect. Indeed, in Dudik et al. (2007) the [Ne V] fluxes were obtained using the short-wavelength, high-resolution (SH; 4.7″ 11.3″, = 9.9-19.6 m) and long-wavelength, high-resolution (LH; 11.1″ 22.3″, = 18.7-37.2 m) modules of IRS. As it is clear from Figure 17, the extension of [Ne V] emission is detected up to 8″ from the nucleus. This means that the SH module surely lost part of the [Ne V] flux, leading to an artificially enhancement of the [Ne V]/[Ne V] line ratio.
@@ Appendix F HOMERUN emission line fluxes
Emission line | Obs. Luminosity | Model | Mod/Obs | AGN frac. | SF frac. | AGN dered. | SF dered. |
(1) | (2) | (3) | (4) | (5) | (6) | (7) | (8) |
H | |||||||
[O III]5007Å | |||||||
[N I]∗5199Å | |||||||
He II 5411Å | |||||||
[Fe VII]5721Å | |||||||
[N II]5755Å | |||||||
[He I]∗5876Å | |||||||
[Fe VII]6087Å | |||||||
[O I]6300Å | |||||||
[S III]6312Å | |||||||
[O I]6364Å | |||||||
[Fe X]6375Å | |||||||
H | |||||||
[N II]6583Å | |||||||
He I 6678Å | |||||||
[S II]6716Å | |||||||
[S II]6731Å | |||||||
[Ar V]7006Å | |||||||
[He I]∗7065Å | |||||||
[Ar III]7136Å | |||||||
Ca II 7291Å | |||||||
[O II]∗7323Å | |||||||
[O II]∗7332Å | |||||||
[Ni II]7378Å | |||||||
[Ar III]7751Å | |||||||
[Cl II]8579Å | |||||||
[Fe II]8617Å | |||||||
[S III]9069Å |
Emission line | Obs. Luminosity | Model | Mod/Obs | AGN frac. | SF frac. | AGN dered. | SF dered. |
(1) | (2) | (3) | (4) | (5) | (6) | (7) | (8) |
[Fe II]5m | |||||||
[Fe VIII]5m | |||||||
[Mg VII]5m | |||||||
[Mg V]5m | |||||||
[Ni II]6m | |||||||
[Fe II]6m | |||||||
[Ar II]7m | |||||||
[Na III]7m | |||||||
Pa | |||||||
[Ne VI]7m | |||||||
[Fe VII]7m | |||||||
[Ar V]8m | |||||||
[Na VI]8m | |||||||
[Ar III]9m | |||||||
[Na IV]9m | |||||||
[Fe VII]9m | |||||||
[S IV]10m | |||||||
Hu | |||||||
[Ne II]13m | |||||||
[Ar V]13m | |||||||
[Mg V]13m | |||||||
[Ne V]14m | |||||||
[Cl II]14m | |||||||
[Na VI]14m | |||||||
[Ne III]16m | |||||||
[Fe II]18m | |||||||
[S III]18m | |||||||
[Fe VI]19m | |||||||
[Fe III]23m | |||||||
[Ne V]24m | |||||||
[O IV]26m | |||||||
[Fe II]26m |
@@ Appendix G Ionized gas mass estimate
We corrected both the [Ne V] and the [O III] luminosity for dust attenuation using the Balmer decrement from the Hα/Hβ ratio, employing the Gordon et al. (2023) attenuation law and an intrinsic ratio Hα/Hβ = 2.86 (for Te = 104 K, Osterbrock & Ferland 2006). Assuming the two different densities for the optical and mid-IR species previously computed, we obtain an outflow mass of = (2.6 1.0) 103 and = (20 7) 104 .
Appendix H Moment maps
In this appendix, we present the moment maps of H2, recombination, low IP, medium IP, and high IP emission lines from Fig. 13 to Fig. 17, respectively.

### v3

@@ C.1 JWST/MIRI MRS emission line fitting
This appendix provides further details on the customized python routine we employed to analyze the MIRI MRS datacubes. We followed a three-step procedure: first, we applied spatial smoothing to enhance the S/N; second, we subtracted the local continuum around the brightest emission lines; finally, we performed spaxel-by-spaxel Gaussian fitting of the residual emission. First, we applied a Gaussian spatial smoothing with a kernel size of 1 pixel to every spatial plane of the datacube. This step improves the S/N of faint extended features, at the cost of a moderate degradation in spatial resolution. Specifically, the effective PSF is broadened by approximately 40 across all wavelengths, due to the fact that the pixel size is comparable to or slightly smaller than the intrinsic PSF at each wavelength. Nevertheless, the resulting resolution remains sufficient for our spatially resolved analysis, as it still allows us to distinguish the relevant morphological and kinematic structures in the data.
@@
and the integrated flux within the FoV of the respective channel, expressed in units of erg s-1 cm-2.
[…]
@@
In the mid-IR, dust attenuation is much lower than at optical wavelengths, but it does vary with wavelength and affect emission lines differently, potentially impacting specific line ratios. Unfortunately, in our data the S/N of the Pf and Hu (i.e. H I (6-5) at 7.46 m and H I (7-6) at 12.37 m, respectively) is too low to derive a reliable extinction map directly from the mid-IR spectra. To test the potential impact of dust attenuation, we applied an extinction correction to the mid-IR emission lines using the G23 (Gordon et al. 2009, 2021, 2023; Fitzpatrick et al. 2019; Decleir et al. 2022) attenuation curve, from the dustextinction package777https://github.com/karllark/dustextinction (Gordon 2024), and adopting the maximum value (3) derived from our best-fit photoionization model (see Section 3.5.3). However, we find that the resulting variations in the mid-IR line ratios are within the errors, including those involving emission lines at distant wavelengths (e.g., [Ne V]/[Ne V], see Section 3.4).
@@ C.2 MUSE WFM emission line fitting
MUSE data were analyzed as already described in previous works of the MAGNUM program (Venturi et al. 2018; Mingozzi et al. 2019). Briefly, after a preliminary Voronoi tessellation (Cappellari & Copin 2003) to achieve an average S/N per wavelength channel of at least 70 on the continuum, we applied the penalized PiXel-Fitting (pPXF, Cappellari & Emsellem 2004) code to measure the optical emission lines listed in Table 3. Using the penalized PiXel-Fitting (pPXF, Cappellari & Emsellem 2004) code, we fitted simultaneously the continuum and gas emission in the wavelength range 4800-9200 Å, convolving a linear combination of stellar templates (E-MILES; for details see Vazdekis et al. 2016) with a Gaussian velocity distribution. Each emission line was fitted twice, using one and two Gaussian components, and the KS test was applied spaxel-by-spaxel to determine the minimum number of components required to adequately reproduce the line profiles (see Appendix C.1). Given that NGC 1365 hosts a Seyfert 1.8 nucleus, we included an additional spectral component when fitting the bins within a radius of 10 spaxels from the nuclear position to account for the spatially unresolved broad-line region (BLR). The BLR model consists of two independent components: a broad (¿1000 km s-1) Gaussian component for the broad Balmer lines (i.e. H and H) and a template of broad [Fe II] emission features by Boroson & Green (1992). The nuclear model spectrum was obtained by fitting the integrated spectrum of the central spaxels and, in the subsequent spaxel-by-spaxel fitting, it was kept fixed in shape and velocity while being allowed to scale in flux. We then subtracted the stellar and nuclear emission from the original unbinned data cube by rescaling the fitted continuum - assumed constant within each Voronoi bin - to match the observed continuum flux in each spaxel, before performing the subtraction.
@@
Finally, we corrected the optical emission-line fluxes for dust attenuation using the Balmer decrement (i.e., H/H), assuming a case-B hydrogen recombination value of 2.86 for this ratio (for an electron temperature and density of Te = K and ne = cm-3, respectively; Osterbrock & Ferland 2006) and adopting the G23 attenuation law. The results, presented in Appendix D, are consistent with previous estimates by Venturi et al. (2018) and Mingozzi et al. (2019). As shown in Fig. 12, the dust attenuation peaks at AV = 3.7 in the circumnuclear ring, where young stellar clusters are embedded (Liu et al. 2023; Schinnerer et al. 2023). High attenuation values are also found at the nuclear position, while we do not find any correlation of AV with the outflow morphology, indicative of its low dust content.
@@ Appendix D Extinction and electron density from MUSE
In this appendix, we present the spatially resolved maps of visual extinction and electron density in the circumnuclear region of NGC 1365, derived from the MUSE data and zoomed-in FoV matching the MIRI one. In the top panel of Fig. 12, we show the visual extinction computed from the Balmer decrement (H/H), adopting the G23 extinction law. High visual extinction values of A 3.7 mag are found at the nuclear position and in the southeastern part of the FoV, corresponding to the circumnuclear ring, as also observed by Venturi et al. (2018) and Mingozzi et al. (2019).
@@ Appendix E Density comparison with IRS observations
In this appendix, we compare our spatially resolved JWST/MIRI density measurements with those obtained from integrated spectra by Dudik et al. (2007). Dudik et al. (2007) used the Infrared Spectrograph (IRS) on board Spitzer to compute the [Ne V] ratio in a sample of local AGN. They found that Type 2 AGN show higher line ratios with respect to Type 1, with all Type 2 sources lying above the theoretical LDL threshold, implying log(ne/cm-3)2. In particular, for NGC 1365 they computed a [Ne V]/[Ne V] ratio of 2.40.6, which is larger than our estimation of 1.00.2 obtained as the ratio between the two integrated fluxes in the MIRI FoV. They argued that the [Ne V] emitting region likely originates inside the nuclear torus and that differential infrared extinction due to dust in the obscuring torus may be responsible for the ratio above the LDL. Exploiting the spatially resolved [Ne V] emission in our data and the density values in the nuclear region, we can rule out this interpretation. However, given the presence of some LDL regions in Fig. 6, it is possible that their finding is the result of spatial integration over a wide slit, thus including such regions in the measurement. A more plausible reason is an observational effect. Indeed, in Dudik et al. (2007) the [Ne V] fluxes were obtained using the short-wavelength, high-resolution (SH; 4.7″ 11.3″, = 9.9-19.6 m) and long-wavelength, high-resolution (LH; 11.1″ 22.3″, = 18.7-37.2 m) modules of IRS. As it is clear from the moment maps of the high-IP species in the online material, the extension of [Ne V] emission is detected up to 8″ from the nucleus. This means that the SH module surely lost part of the [Ne V] flux, leading to an artificially enhancement of the [Ne V]/[Ne V] line ratio.
@@ Appendix F HOMERUN emission line fluxes
| Emission line | [Å] | Obs. Luminosity | Model | Mod/Obs | AGN frac. | SF frac. | AGN dered. | SF dered. |
| (1) | (2) | (3) | (4) | (5) | (6) | (7) | (8) | (9) |
| H | 4861 | |||||||
| [O III] | 5007 | |||||||
| [N I]∗ | 5199 | |||||||
| He II | 5411 | |||||||
| [Fe VII] | 5721 | |||||||
| [N II] | 5755 | |||||||
| [He I]∗ | 5876 | |||||||
| [Fe VII] | 6087 | |||||||
| [O I] | 6300 | |||||||
| [S III] | 6312 | |||||||
| [O I] | 6364 | |||||||
| [Fe X] | 6375 | |||||||
| H | 6563 | |||||||
| [N II] | 6583 | |||||||
| He I | 6678 | |||||||
| [S II] | 6716 | |||||||
| [S II] | 6731 | |||||||
| [Ar V] | 7006 | |||||||
| [He I]∗ | 7065 | |||||||
| [Ar III] | 7136 | |||||||
| Ca II | 7291 | |||||||
| [O II]∗ | 7323 | |||||||
| [O II]∗ | 7332 | |||||||
| [Ni II] | 7378 | |||||||
| [Ar III] | 7751 | |||||||
| [Cl II] | 8579 | |||||||
| [Fe II] | 8617 | |||||||
| [S III] | 9069 |
| Emission line | [m] | Obs. Luminosity | Model | Mod/Obs | AGN frac. | SF frac. | AGN dered. | SF dered. |
| (1) | (2) | (3) | (4) | (5) | (6) | (7) | (8) | (9) |
| [Fe II] | 5.34 | |||||||
| [Fe VIII] | 5.44 | |||||||
| [Mg VII] | 5.50 | |||||||
| [Mg V] | 5.61 | |||||||
| [Ni II] | 6.64 | |||||||
| [Fe II] | 6.72 | |||||||
| [Ar II] | 6.98 | |||||||
| [Na III] | 7.32 | |||||||
| Pa | 7.46 | |||||||
| [Ne VI] | 7.65 | |||||||
| [Fe VII] | 7.81 | |||||||
| [Ar V] | 7.90 | |||||||
| [Na VI] | 8.61 | |||||||
| [Ar III] | 8.99 | |||||||
| [Na IV] | 9.04 | |||||||
| [Fe VII] | 9.53 | |||||||
| [S IV] | 10.51 | |||||||
| Hu | 12.37 | |||||||
| [Ne II] | 12.81 | |||||||
| [Ar V] | 13.10 | |||||||
| [Mg V] | 13.52 | |||||||
| [Ne V] | 14.32 | |||||||
| [Cl II] | 14.37 | |||||||
| [Na VI] | 14.40 | |||||||
| [Ne III] | 15.56 | |||||||
| [Fe II] | 17.94 | |||||||
| [S III] | 18.71 | |||||||
| [Fe VI] | 19.56 | |||||||
| [Fe III] | 22.93 | |||||||
| [Ne V] | 24.32 | |||||||
| [O IV] | 25.89 | |||||||
| [Fe II] | 25.99 |
@@ Appendix G Ionized gas mass estimate
We corrected both the [Ne V] and the [O III] luminosity for dust attenuation using the Balmer decrement from the Hα/Hβ ratio, employing the G23 attenuation law and an intrinsic ratio Hα/Hβ = 2.86 (for Te = 104 K, Osterbrock & Ferland 2006). Assuming the two different densities for the optical and mid-IR species previously computed, we obtain an outflow mass of = (2.6 1.0) 103 and = (20 7) 104 .